OpenStudy (anonymous):

x - 7(squrare root of x) =-12

OpenStudy (anonymous):

\[x-7\sqrt{x}=-12\]

OpenStudy (anonymous):

the answer is 9

OpenStudy (anonymous):

it has two answers its either 9 and 4 or 9 and 11

OpenStudy (anonymous):

not 11 i mean 16

OpenStudy (anonymous):

9 and 16

OpenStudy (anonymous):

\[x-7\sqrt{x}=-12\] \[x+12=7\sqrt{x}\] \[x^{2}+24x+144=49x\] \[x^{2}-25x+144=0\] Solve for x.

OpenStudy (anonymous):

ohk so hold on i got another one

OpenStudy (anonymous):

\[x - 5\sqrt{x} +6=0\]

OpenStudy (anonymous):

@AnanyaManas Please don't just give out the answer w/ no work. It's against the code of conduct. @janessia Please start a new question and close this one.

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

i was typing the answer but then i saw that you have alredy given one. @Wired

OpenStudy (anonymous):

@Wired where did you get 24 from?

OpenStudy (anonymous):

\[x+12=7\sqrt{x}\] Square both sides: \[(x+12)^2=7(\sqrt{x})^{2}\]